OpenStudy (anonymous):

logbase3(x)+logbase3(x+6)=3

OpenStudy (anonymous):

\[\log_3(x)+\log_3(x+6)=\log_3(x(x+6))\] is a start

OpenStudy (anonymous):

\[\log_{3} (x^2+6x)=3\] then i get lost

OpenStudy (anonymous):

rewrite in equivalent exponential form \[\log_b(x)=y\iff b^y=x\]

OpenStudy (anonymous):

you should be able to switch back and forth easily will make life much easier math life anyways

OpenStudy (anonymous):

if i do \[\log_{3} 3 \] then that equals one, and x^2+6x=1 I can't factor if I bring it to the other side, the 3 becomes a 27 somehow. What property is it called

OpenStudy (anonymous):

\[\log_7(49)=2\iff 7^2=49\] etc

OpenStudy (anonymous):

it is the meaning of the log

OpenStudy (anonymous):

\[10^3=1000\iff \log_{10}(1000)=3\]

OpenStudy (anonymous):

Oh I see... but logbase3(3) = 1

OpenStudy (anonymous):

the teacher got 27 somehow, as 3^1 = 3

OpenStudy (anonymous):

lets go slow you are not "taking the log" of both sides

OpenStudy (anonymous):

Ok, thank you for your patience first of all.... I am combining logs using a property on the left side

OpenStudy (anonymous):

how would you solve \[\log_2(x)=3\] for \(x\) ?

OpenStudy (anonymous):

2^3

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

how would you solve \[\log_3(x)=3\]?

OpenStudy (anonymous):

3^3.. haha

OpenStudy (anonymous):

found that 27 after all didn't you?

OpenStudy (anonymous):

so the reason I give the 3 on the right side logbase3 is to make my life easier to work on the left side, correct?

OpenStudy (anonymous):

:) yes.

OpenStudy (anonymous):

you do not "give" anything \[\log_3(\text{whatever})=3\iff \text{whatever}=3^3\]

OpenStudy (anonymous):

oh ok, i understand... thank you so much